# Each of those eight coalition logics is also determined by six other kinds of models

### v1

@@ 5.3 Representation of single-coalition-first action models by single-coalition-first neighborhood models
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood $\mathtt{X}$-model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
@@
  • $\mathtt{AC}=\{\alpha_{a-s-X}\mid a\in\mathtt{AG},s\in\mathtt{ST},\text{ and }X\in\mathtt{nei}_{a}(s)\}$.
  • for every $a\in\mathtt{AG}$, $s\in\mathtt{ST}$ and $\alpha_{x-u-X}\in\mathtt{AC}$:
    $\mathsf{out}_{a}(s,\alpha_{x-u-X})=\begin{cases}X&\text{if $a=x$ and $u=s$}\\ \emptyset&\text{otherwise}\end{cases}$
@@
Let $\{\mathtt{AE}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of actual effectivity functions of $\mathtt{S}\textbf{-}\mathtt{AM}$. Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$. It suffices to show that $\mathtt{AE}_{\mathrm{C}}(s)=\mathtt{nei}_{\mathrm{C}}(s)$.
Assume $\mathrm{C}=\emptyset$.
Then, $\mathtt{AE}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}=\mathtt{nei}_{\mathrm{C}}(s)$.
Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$.
Note:
$\mathtt{nei}_{\mathrm{C}}(s)=\bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}$
$\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{\mathrm{C}}(s,\sigma_{a_{n}})$
$\mathtt{AE}_{\mathrm{C}}(s)=\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)\}$
Let $X\in\mathtt{AE}(s,\mathrm{C})$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that $X=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Let $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$, where $\sigma_{a_{1}}\in\mathtt{JA}_{a_{1}},\dots,\sigma_{a_{n}}\in\mathtt{JA}_{a_{n}}$. Note that $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}(s,\sigma_{a_{n}})$ and $X$ is not empty. Then, $\sigma_{a_{1}}\in\mathtt{av}_{a_{1}}(s),\dots,\sigma_{a_{n}}\in\mathtt{av}_{a_{n}}(s)$.
Let $1\leq i\leq n$ and $\sigma_{a_{i}}=(a_{i},\alpha_{x-u-X_{i}})$. Note $\mathsf{out}_{a_{i}}(s,\sigma_{a_{i}})\neq\emptyset$. Then, $\mathsf{out}(s,\sigma_{a_{i}})=X_{i}$, $\sigma_{a_{i}}=(a_{i},\alpha_{a_{i}-s-X_{i}})$ and $X_{i}\in\mathtt{nei}_{a_{i}}(s)$. Then, $X_{1}\cap\dots\cap X_{n}=X$. Then, $X\in\mathtt{nei}(s,\mathrm{C})$.
Let $X\in\mathtt{nei}(s,\mathrm{C})$. Then, there are $X_{1}\in\mathtt{nei}_{a_{1}}(s),\dots,X_{n}\in\mathtt{nei}_{a_{n}}(s)$ such that $X_{1}\cap\dots\cap X_{n}=X$. Then, $\alpha_{a_{1}-s-X_{1}}\in\mathtt{AC},\dots,\alpha_{a_{n}-s-X_{n}}\in\mathtt{AC}$. Let $\sigma_{a_{1}}=\{(a_{1},\alpha_{a_{1}-s-X_{1}})\},\dots,\sigma_{a_{n}}=\{(a_{n},\alpha_{a_{n}-s-X_{n}})\}$. Then, $\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})=X_{1},\dots,\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})=X_{n}$. Let $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=X_{1}\cap\dots\cap X_{n}=X$. Note $X$ is not empty. Then, $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)$. Then, $X\in\mathtt{AE}_{\mathrm{C}}(s)$.
@@
For every single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}$ and single-coalition-first neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}$, if $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by $\mathtt{S}\textbf{-}\mathtt{NM}$, then for all $\mathtt{X}\in\mathrm{ES}$, $\mathtt{S}\textbf{-}\mathtt{AM}$ is an $\mathtt{X}$-model if and only if $\mathtt{S}\textbf{-}\mathtt{NM}$ is an $\mathtt{X}$-model.
@@
Let $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathtt{AE}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of actual effectivity functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
Assume $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by $\mathtt{S}\textbf{-}\mathtt{NM}$. Then for all $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{AE}_{\mathrm{C}}=\mathtt{nei}_{\mathrm{C}}$.
First, it is easy to see the following statements are equivalent:
@@
Second, it is easy to check the following statements are equivalent:
@@
  • for all $s\in\mathtt{ST},\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset,\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}$, if $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$, then $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$;
@@
Then, it follows that for all $\mathtt{X}\in\mathrm{ES}$, $\mathtt{S}\textbf{-}\mathtt{AM}$ is an $\mathtt{X}$-model if and only if $\mathtt{S}\textbf{-}\mathtt{NM}$ is an $\mathtt{X}$-model.
Theorem 8 (Representation of single-coalition-first action models by single-coalition-first neighborhood models).
@@
Proof.
Let $\mathtt{X}\in\mathrm{ES}$.
Let $\mathtt{S}\textbf{-}\mathtt{AM}$ be a single-coalition-first action $\mathtt{X}$-model. By Theorem 5, $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by a single-coalition-first neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}$. By Theorem 7, $\mathtt{S}\textbf{-}\mathtt{NM}$ is a single-coalition-first neighborhood $\mathtt{X}$-model.
Let $\mathtt{S}\textbf{-}\mathtt{NM}$ be a single-coalition-first neighborhood $\mathtt{X}$-model. By Theorem 6, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}$. By Theorem 7, $\mathtt{S}\textbf{-}\mathtt{AM}$ is a single-coalition-first action $\mathtt{X}$-model.
@@ 6.1 Clear grand-coalition-first action models
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, and let $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$.
@@
    for all $a\in\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$, if $\sigma_{a}\neq\sigma^{\prime}_{a}$, then $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})=\emptyset$;
@@
    for all $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, if $\sigma_{\mathrm{C}}\neq\sigma^{\prime}_{\mathrm{C}}$, then $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})=\emptyset$;
@@
    for all $s\in\mathtt{ST}$, and $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, if $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$, then $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$.
@@
Assume (1). Assume (2) does not hold. Then, there is $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ such that $\sigma_{\mathrm{C}}\neq\sigma^{\prime}_{\mathrm{C}}$ and $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})\neq\emptyset$.
Let $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})$. Then, there is $\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$, $\sigma^{\prime}_{\mathrm{C}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$, $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$ and $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$.
Then, there is $a\in\mathrm{C},\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$ such that $\sigma_{a}\subseteq\sigma_{\mathrm{C}}$, $\sigma^{\prime}_{a}\subseteq\sigma^{\prime}_{\mathrm{C}}$ and $\sigma_{a}\neq\sigma^{\prime}_{a}$. Then, $\sigma_{a}\subseteq\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{a}\subseteq\sigma^{\prime}_{\mathtt{AG}}$.
Then, $w\in\mathsf{out}_{a}(s,\sigma_{a})$ and $w\in\mathsf{out}_{a}(s,\sigma^{\prime}_{a})$. Then, $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})\neq\emptyset$. There is a contradiction.

### v2

@@ 4.4 Representation of single-coalition-first action models by single-coalition-first neighborhood models
  • $\mathtt{AC}=\{\alpha_{a-s-\mathrm{X}}\mid a\in\mathtt{AG},s\in\mathtt{ST},\text{and }\mathrm{X}\in\mathtt{nei}_{a}(s)\}$.
  • for every $a\in\mathtt{AG}$, $s\in\mathtt{ST}$ and $\alpha_{x-u-\mathrm{X}}\in\mathtt{AC}$,
    $\mathsf{out}_{a}(s,\alpha_{x-u-\mathrm{X}})=\begin{cases}\mathrm{X}&\text{if $x=a$ and $u=s$}\\ \emptyset&\text{otherwise}\end{cases}$
@@
Let $\{\mathtt{AE}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of actual effectivity functions of $\mathtt{S}\textbf{-}\mathtt{AM}$. Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$. It suffices to show $\mathtt{AE}_{\mathrm{C}}(s)=\mathtt{nei}_{\mathrm{C}}(s)$. Note $\mathtt{AE}_{\mathrm{C}}(s)=\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\text{ and }\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\}$.
Assume $\mathtt{suc}(s)=\emptyset$. Note $\mathtt{AE}_{\mathrm{C}}(s)=\emptyset$ and $\mathtt{nei}_{\mathrm{C}}(s)=\emptyset$. Then, $\mathtt{AE}_{\mathrm{C}}(s)=\mathtt{nei}_{\mathrm{C}}(s)$.
Assume $\mathtt{suc}(s)\neq\emptyset$ and $\mathrm{C}=\emptyset$. Note $\mathtt{AE}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}$ and $\mathtt{nei}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}$. Then, $\mathtt{AE}_{\mathrm{C}}(s)=\mathtt{nei}_{\mathrm{C}}(s)$.
Assume $\mathtt{suc}(s)\neq\emptyset$ and $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$. Note $\mathtt{nei}_{\mathrm{C}}(s)=\bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}$.
Let $\mathrm{X}\in\mathtt{AE}_{\mathrm{C}}(s)$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that $\mathrm{X}=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Let $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$. Note that $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{1}}(s,\sigma_{a_{n}})$ and $\mathrm{X}\neq\emptyset$. Then, $\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\neq\emptyset$, …, $\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})\neq\emptyset$. Let $i$ be such that $1\leq i\leq n$ and $\sigma_{a_{i}}=\alpha_{x-u-\mathrm{X}_{i}}$. By the definition of $\mathsf{out}_{a_{i}}(s,\sigma_{a_{i}})$, $\mathsf{out}_{a_{i}}(s,\sigma_{a_{i}})=\mathrm{X}_{i}$, $x=a_{i}$, and $u=s$. Then $\sigma_{a_{i}}=\alpha_{a_{i}-s-\mathrm{X}_{i}}$ and $\mathrm{X}_{i}\in\mathtt{nei}_{a_{i}}(s)$. Then, $\mathrm{X}=\mathrm{X}_{1}\cap\dots\cap\mathrm{X}_{n}$. Then, $\mathrm{X}\in\mathtt{nei}_{\mathrm{C}}(s,\mathrm{C})$.
Let $\mathrm{X}\in\mathtt{nei}_{\mathrm{C}}(s,\mathrm{C})$. Then, $\mathrm{X}=\mathrm{X}_{1}\cap\dots\cap\mathrm{X}_{n}$ for some $\mathrm{X}_{1}\in\mathtt{nei}_{a_{1}}(s),\dots,\mathrm{X}_{n}\in\mathtt{nei}_{a_{n}}(s)$. Then, $\alpha_{a_{1}-s-\mathrm{X}_{1}}\dots,\alpha_{a_{n}-s-\mathrm{X}_{n}}\in\mathtt{AC}$. Note $\mathsf{out}_{a_{1}}(s,\alpha_{a_{1}-s-\mathrm{X}_{1}})=\mathrm{X}_{1}$, $\dots$, $\mathsf{out}_{a_{n}}(s,\alpha_{a_{n}-s-\mathrm{X}_{n}})=\mathrm{X}_{n}$. Let $\sigma_{\mathrm{C}}=\alpha_{a_{1}-s-\mathrm{X}_{1}}\cup\dots\cup\alpha_{a_{n}-s-\mathrm{X}_{n}}$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\alpha_{a_{1}-s-\mathrm{X}_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\alpha_{a_{n}-s-\mathrm{X}_{n}})=\mathrm{X}_{1}\cap\dots\cap\mathrm{X}_{n}=\mathrm{X}$. Note $\mathrm{X}$ is not empty. Then, $\mathrm{X}\in\mathtt{AE}_{\mathrm{C}}(s)$.
@@
For every single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}$ and single-coalition-first neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}$, if $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by $\mathtt{S}\textbf{-}\mathtt{NM}$, then for every $\mathtt{X}\in\mathrm{ES}$, $\mathtt{S}\textbf{-}\mathtt{AM}$ is an $\mathtt{X}$-model if and only if $\mathtt{S}\textbf{-}\mathtt{NM}$ is an $\mathtt{X}$-model.
@@
Let $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathtt{AE}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of actual effectivity functions of $\mathtt{S}\textbf{-}\mathtt{AM}$. Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$. Assume $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by $\mathtt{S}\textbf{-}\mathtt{NM}$. Then, for all $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{AE}_{\mathrm{C}}=\mathtt{nei}_{\mathrm{C}}$.
Let $\mathtt{X}\in\mathrm{ES}$. It is easy to see that the following three groups of statements are equivalent, respectively.
  1. (1)
@@
  2. (2)
@@
    • for all $s\in\mathtt{ST},\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset,\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}$, if $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$, then $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$;
@@
It follows that $\mathtt{S}\textbf{-}\mathtt{AM}$ is an $\mathtt{X}$-model if and only if $\mathtt{S}\textbf{-}\mathtt{NM}$ is an $\mathtt{X}$-model.
The following result follows from Theorems 5, 6, and 7.
Theorem 8 (Representation of the class of single-coalition-first action $\mathtt{X}$-models by the class of single-coalition-first neighborhood $\mathtt{X}$-models).
@@ 5.1 Clear grand-coalition-first action models
The following result offers two alternative definitions of clear grand-coalition-first action models.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, $\mathtt{suc}$ be the sucessor function of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $s\in\mathtt{ST}$.
@@
    for all $a\in\mathtt{AG}$ and $\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$, if $\sigma_{a}\neq\sigma^{\prime}_{a}$, then $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})=\emptyset$;
@@
    for all $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, if $\sigma_{\mathrm{C}}\neq\sigma^{\prime}_{\mathrm{C}}$, then $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})=\emptyset$;
@@
    for all $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, if $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$, then $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$.
@@
Assume (1). Suppose (2) does not hold. Then, there is $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ such that $\sigma_{\mathrm{C}}\neq\sigma^{\prime}_{\mathrm{C}}$ and $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})\neq\emptyset$. Then, there is $a\in\mathrm{C}$ and $\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$ such that $\sigma_{a}\subseteq\sigma_{\mathrm{C}}$, $\sigma^{\prime}_{a}\subseteq\sigma^{\prime}_{\mathrm{C}}$, and $\sigma_{a}\neq\sigma^{\prime}_{a}$. Let $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})$. Then, there is $\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$, $\sigma^{\prime}_{\mathrm{C}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$, $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$, and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$. Then, $\sigma_{a}\subseteq\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{a}\subseteq\sigma^{\prime}_{\mathtt{AG}}$. Then, $t\in\mathsf{out}_{a}(s,\sigma_{a})$ and $t\in\mathsf{out}_{a}(s,\sigma^{\prime}_{a})$. Then, $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})\neq\emptyset$, which is impossible.